The sureface are of a cube is 384 sq. inches. FInd the lenght of an edge.
A. 24
B. 4
c. 8
D. 16

Answer by stanbon(75887)

The surface area of a cube is 384 sq. inches. FInd the length of an edge.
----
The cube has 6 faces.
Each face has area = e^2
Surface area = 6*e^2 where e is the length of one edge.
-----
Solve:
6e^2 = 384
e^2 = 64
edge = 8 inches
---------------------
Cheers,
Stan H.
====================
A. 24
B. 4
c. 8
D. 16